OpenStudy (anonymous):

What's the inverse of: ln(2x)-3

OpenStudy (bradely):

y =ln(2x)-3 Add 3 both sides y+3 =ln(2x) 2x =e^(y + 3) x =[e^(y + 3)]/2 Replace f^(-1) (x) =[e^(x + 3)]/2 Source: http://www.mathskey.com/question2answer/